A square has a perimeter of 58 cms. What is the area of the square? please show work thanks so much :D
--sandy

Answer by checkley71(8403)

4X=58 OR X=58/4 OR X=14.5 CMS.FOR THE SIDES THEREFORE THE AREA=14.5^2 OR AREA=210.25 SQ CMS.